The equation of displacement of particle performing SHM is X = 0.25 sin (200 t). The maximum velocity is
(a) 100 m/s (b) 200 m/s
(c) 50 m/s (d) 150 m/s

Grade:Upto college level

1 Answers

Vikas TU
3 years ago
v = dx /dt 
Maximum amplitude is maximum velocity 
v = 0.25 * 200 = 50 m/s 
C is correct 
….................................